# The Open mulTiwavelength Transient Event Repository (OTTER): Infrastructure Release and Tidal Disruption Event Catalog

### v1

@@ 5.4 Example 4: Validating Optical Transient Classifiers in the Rubin era
A sample application of the OTTER dataset is to apply a machine learning classifier to the available optical data to further test the ability of the classifier to recover the TDE classification from a sample of spectroscopically classified TDEs. We use the Finding Luminous and Exotic Extragalactic Transients (FLEET) algorithm, which is able to classify transients based on their optical light curves as SLSN, TDE, and “other” (Gomez et al., 2020, 2023) [^22] Note, while we report other classifications here, the classifications besides TDE and SLSN are not validated.. FLEET requires at least 4 detections in $g$-band or $r$-band, and is able to work with SDSS/ZTF filters (i.e. ugrizy) so we select all transients in OTTER that have enough photometry in these filters to model and run FLEET. This results in 83 transients, all of which are spectroscopically classified as TDEs.
@@
A final notable result is that FLEET appears to fail on the two TDEs in this sample that show on-axis jets (Sw J1644+57 and AT 2022cmc), returning P(TDE)=0% and finding very low probabilities for all classifications that FLEET is trained on. This is unsurprising given that these TDEs tend to appear much different photometrically from the non-jetted population (although they may share some similarities with the most luminous optical TDEs; Yao et al. 2025). Additionally, given the extremely limited sample size of jetted TDEs, FLEET did not include any in their training sample. This may indicate that future TDE classifiers should attempt to classify jetted and non-jetted TDEs separately. In the future, we hope that OTTER will become a hub for curating training and validation datasets for future machine learning classifiers, particularly those that utilize multiwavelength datasets.
6 Conclusions
@@
For the initial release of this infrastructure, we also present the largest catalog of tidal disruption events (TDEs). As of September 2025, the catalog includes $\sim 82,000$ UVOIR, $\sim 1,200$ X-ray, and $\sim 2,300$ radio photometry points from 232 TDEs that were classified in at least one published source. The API and web-application allow for easy access to this dataset, and the OTTER upload form will allow others to upload their datasets after publishing so that, as a field, we can continue to maintain this catalog.
We present four sample analyses using the available TDE data including a delay time analysis, a sub-population analysis, simple modeling, and validating a machine learning classifier. First, we perform a simple analysis of the TDE population and compute the delay time between the peak radio and X-ray emission to be $\sim 20$ days. Second, we briefly discuss an analysis of the population of known Extreme Coronal Line Emitters (ECLEs) in the context of the broader TDE population. We recover the apparent relation that ECLEs appear to occur in TDE hosts with more gas rich circumnuclear environments with a high statistical significance. Third, we model two TDEs with MOSFiT to show that it is very easy to apply models to this standard dataset. Fourth, and finally, we apply the FLEET transient classifier (Gomez et al., 2020, 2023) to the data in OTTER and find that FLEET successfully classifies $\sim 69\%$ of the sample as TDEs.
@@
We thank Edo Berger for useful discussion and feedback, particularly on the classification confidence flagging. We thank the staff at SciServer, namely Charles Glaser, Gerard Lemson, and Arik Mitschang, for their indispendable discussion and assistance deploying OTTER and making it accessible to the world. We thank Ryan Chornock, Iair Arcavi, Sara Faris, Margaret Shepherd, and Yvette Cendes for testing the OTTER web interface and providing feedback. We thank Megan Masterson and Muryel Guolo for taking the time to provide data from their respective population papers. We thank Mathieu Renzo for useful comments on the paper draft. N.F. acknowledges support from the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-2137419. KDA and CTC acknowledge support provided by the NSF through award SOSPA9-007 from the NRAO and award AST-2307668. KDA gratefully acknowledges support from the Alfred P. Sloan Foundation. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
Appendix A Metadata and References for each TDE
@@
The default name, coordinates, redshift and discovery date (i.e. the metadata) for all of the TDEs in OTTER is given in Table LABEL:tab:refs. The rightmost column gives all of the references associated with this transient event in OTTER, including metadata and photometry. The reference numbers are defined below LABEL:tab:refs.
Table 3: Default metadata information for the transients in OTTER. C is the confidence flag associated with the default classification of the transient.
Transient Name RA [HMS] Dec [DMS] Redshift Discovery Date C References
RX J1420+53 14:20:24.2 +53:34:11 0.147 1990-07-11 1.0 4 , 74 , 98 , 100 , 346
NGC 5905 15:15:23.32 +55:31:01.59 0.01124 1990-07-11 3.1 1 , 6 , 43 , 74 , 98 , 100 , 202 , 346
RX J1624+75 16:24:56.7 +75:54:57.5 0.0636 1990-08-25 1.0 3 , 43 , 74 , 98 , 100 , 202 , 346
RBS 1032 11:47:26.8 +49:42:59 0.026 1990-11-05 1.0 54 , 74 , 98 , 100 , 346
LEDA 095953 13:47:30.33 -32:54:50.63 0.0366 1991-07-06 1.0 11 , 74 , 98 , 100 , 346
RX J1242-11A 12:42:38.55 -11:19:20.8 0.05 1992-07-15 1.0 2 , 4 , 43 , 74 , 98 , 100 , 202 , 346
FIRST J153350.8+272729 15:33:50.88 +27:27:29.618 0.032430 1995-04-16 1.0 286
WINGS J1348 13:48:49.9 +26:35:57.49 0.062 1999-12-20 1.0 47 , 64 , 74 , 98 , 100 , 346
SDSSJ0159 01:59:57.64 +00:33:10.49 0.31167 2000-07-29 0.0 62 , 68 , 74 , 98 , 100 , 346
3XMM J152130.7+074916 15:21:30.75 +07:49:16.7 0.17901 2000-08-23 1.0 66 , 74 , 98 , 100 , 246 , 346
NGC 3599 11:15:26.95 +18:06:37.33 0.002699 2002-05-27 1.0 8 , 74 , 98 , 100 , 106 , 346
CSS100217 10:29:13 +40:42:20 0.147 2002-12-31 3.1 5 , 14 , 22 , 27 , 74 , 98 , 100 , 103 , 114 , 346
IGR J17361-4441 17:36:17.42 -44:44:05.98 0.000271 2003-03-22 1.0 17 , 56 , 74 , 98 , 100 , 346
2XMM J123103.2+110648 12:31:03.2 +11:06:48 0.11871 2003-07-13 0.0 36 , 37 , 46 , 74 , 98 , 110 , 346
SDSSJ0748 07:48:20.67 +47:12:14.23 0.0615 2003-10-23 3.1 25 , 34 , 74 , 98 , 100 , 346
SDSSJ1323 13:23:41.97 +48:27:01.26 0.08754 2003-12-02 1.0 8 , 43 , 74 , 98 , 100 , 106 , 202 , 346
D3-13 14:19:29.78806608 +52:52:06.3714576 0.3698 2004-01-15 3.3 7 , 9 , 21 , 42 , 74 , 98 , 100 , 111 , 202 , 346 , 379
SDSSJ1311 13:11:22.2 -01:23:45.2 0.18 2004-02-28 3.3 16 , 43 , 74 , 98 , 100 , 202 , 346
D1-9 02:25:16.96991014 -04:32:59.13393086 0.326 2004-06-22 3.3 7 , 9 , 42 , 74 , 98 , 100 , 202 , 346 , 379
2MASX J0249 02:49:17.28 -04:12:54 0.0186 2004-07-24 1.0 8 , 73 , 74 , 98 , 100 , 346
NGC 6021 15:57:30.72 +15:57:21.6 0.015607 2005-01-02 0.0 71 , 74 , 98 , 100 , 346
UGC 01791 02:19:53.52 +28:14:52.8 0.015881 2005-01-09 1.0 71 , 74 , 98 , 100 , 346
@@
PGC 133344 21:42:55.92 -30:07:58.8 0.02365 2005-03-17 1.0 71 , 74 , 98 , 100 , 346
3XMM J215022.4-055108 21:50:22 -05:51:08 0.055 2005-05-01 1.0 131 , 346
XMMSL1 J131952.3+225958 13:19:52 +22:59:58 0.023 2005-07-15 1.0 191 , 346
3XMM J150052.0+015452 15:00:52 +01:54:52 0.145 2005-07-23 1.0 74 , 98 , 115 , 346
PGC 170392 22:26:46.32 -15:01:22.8 0.016246 2005-11-28 1.0 71 , 74 , 98 , 100 , 346
XMMSL2 J144605.0+685735 14:46:05 +68:57:31 0.029 2006-08-22 3.3 150 , 202 , 346
TDE1 23:42:01.406 +01:06:29.3 0.136 2006-08-27 3.3 5 , 26 , 42 , 74 , 98 , 100 , 111 , 202 , 346 , 379
2XMMi J1847 18:47:25.16 -63:17:24.96 0.0353 2006-09-01 1.0 24 , 74 , 98 , 127 , 346
PGC 1127938 01:18:56.64 -01:03:10.8 0.02 2007-01-08 1.0 71 , 74 , 98 , 100 , 346
TDE2 23:23:48.615 -01:08:10.34 0.2515 2007-06-18 3.3 5 , 26 , 28 , 42 , 74 , 98 , 100 , 202 , 346 , 379
D23H-1 23:31:59.5344 +00:17:14.568 0.1855 2007-09-04 3.3 12 , 42 , 74 , 98 , 100 , 111 , 202 , 346 , 379
SDSSJ0952 09:52:09.56 +21:43:13.24 0.0789 2008-02-05 3.1 10 , 74 , 98 , 100 , 346
Dougie 12:08:47.87 +43:01:20.01 0.191 2009-01-17 3.1 61 , 74 , 98 , 100 , 346
J091225 09:12:25.01 +06:10:14.8 0.1453 2009-02-09 1.0 74 , 98 , 129 , 346
J100933 10:09:33.14 +23:22:55.8 0.0719 2009-02-20 1.0 74 , 98 , 129 , 346
J113527 11:35:27.29 +39:28:10.7 0.1084 2009-04-07 1.0 74 , 98 , 129 , 346
J133737 13:37:37.13 +20:23:51.7 0.0725 2009-04-12 1.0 74 , 98 , 129 , 346
J130819 13:08:19.12 +43:45:25.6 0.0365 2009-04-14 1.0 74 , 98 , 129 , 346
J141036 14:10:36.81 +26:54:25 0.1071 2009-04-24 1.0 74 , 98 , 129 , 346
J142401 14:24:01.66 +29:50:28.4 0.0855 2009-04-26 1.0 74 , 98 , 129 , 346
J145851 14:58:51.65 +17:50:57.9 0.1165 2009-04-29 1.0 74 , 98 , 129 , 346
J121116 12:11:16.47 +00:13:45.3 0.0763 2009-05-04 1.0 74 , 98 , 129 , 346
PTF09ge 14:57:03.1 +49:36:40.8 0.064 2009-05-07 3.3 5 , 13 , 41 , 55 , 74 , 98 , 100 , 111 , 346 , 379
SDSSJ1342 13:42:44.41 +05:30:56.2 0.0366 2009-05-17 1.0 34 , 45 , 74 , 98 , 100 , 129 , 346
J155223 15:52:23.29 +32:34:55.1 0.1277 2009-05-31 1.0 74 , 98 , 129 , 346
PTF09axc 14:53:13.06 +22:14:32.2 0.1146 2009-06-20 3.3 5 , 15 , 41 , 55 , 74 , 98 , 100 , 111 , 202 , 346 , 379
PTF09djl 16:33:55.94 +30:14:16.3 0.184 2009-07-24 3.3 5 , 15 , 41 , 55 , 74 , 98 , 100 , 111 , 113 , 346 , 379
J133837 13:38:37.6 +57:31:13.2 0.1269 2009-11-02 1.0 74 , 98 , 129 , 346
J030257 03:02:57.14 -08:29:42.9 0.1056 2009-11-08 1.0 74 , 98 , 129 , 346
PGC 1190358 15:05:28.56 +01:17:31.2 0.00766 2009-12-29 1.0 71 , 74 , 98 , 100 , 346
J123715 12:37:15.1 +60:12:07 0.2164 2010-01-23 1.0 74 , 98 , 129 , 346
PGC 1185375 15:03:50.4 +01:07:37.2 0.00523 2010-02-03 1.0 71 , 74 , 98 , 100 , 346
PGC 015259 04:29:21.84 -04:45:36 0.014665 2010-02-08 1.0 71 , 74 , 98 , 100 , 346
PS1-10jh 16:09:28.27612107 +53:40:23.99117015 0.1696 2010-05-16 3.3 5 , 40 , 42 , 67 , 74 , 98 , 100 , 111 , 202 , 346 , 379
SDSSJ1201 12:01:36.03 +30:03:05.52 0.146 2010-06-10 3.3 32 , 74 , 98 , 100 , 106 , 202 , 346
PTF10iya 14:38:41 +37:39:33.6 0.22405 2010-06-11 3.3 5 , 38 , 42 , 74 , 98 , 100 , 202 , 346
PTF10nuj 16:26:24.7 +54:42:21.6 0.132 2010-06-13 3.3 5 , 55 , 74 , 98 , 100 , 346
2022agi 01:02:50.007 -22:21:57.22 0.117835 2010-06-25 3.1 74 , 98 , 116 , 319 , 346
2MASX J01190869-3411305 01:19:08 -34:11:30 0.018 2010-07-14 1.0 162 , 346
2021zlg 20:42:44.74 +15:30:32.1 0.203 2010-08-15 3.3 74 , 98 , 114 , 281 , 346
UGC 03317 05:33:37.68 +73:43:26.4 0.004136 2010-09-22 1.0 71 , 74 , 98 , 100 , 346
PS1-11af 09:57:26.82 +03:14:00.924 0.4046 2010-12-30 3.3 5 , 52 , 74 , 98 , 100 , 202 , 346 , 379
IGR J12580+0134 12:58:05.1 +01:34:26 0.00411 2011-01-11 1.0 65 , 80 , 102 , 285
Sw J1644+57 16:44:49.9313 +57:34:59.6893 0.354 2011-03-28 3.3 29 , 30 , 31 , 33 , 39 , 44 , 74 , 76 , 81 , 98 , 100 , 126 , 234 , 346
Swift J2058.4+0516 20:58:19.9 +05:13:32 1.1853 2011-05-27 3.1 18 , 35 , 38 , 63 , 74 , 98 , 100 , 112 , 346
PTF11glr 16:54:06.13 +41:20:14.8 0.207 2011-05-28 3.3 5 , 55 , 74 , 98 , 100 , 346
XMMSL1 J063045.9-603110 06:30:46 -60:31:10 0.014 2011-12-01 1.0 20 , 72 , 93 , 346
Sw J1112-82 11:11:47.6 -82:38:44.44 0.89 2011-12-27 3.1 19 , 69 , 74 , 98 , 100 , 112 , 346
PS1-12yp 13:31:55.91 +23:54:05.7 0.581 2012-04-09 3.3 82 , 98 , 100 , 346
XMMSL1 J061927.1-655311 06:19:27 -65:53:07 0.0729 2012-11-12 3.1 48 , 346
PS1-13jw 08:44:53.56 +42:57:44.8 0.345 2013-02-06 3.3 74 , 98 , 114 , 346
ASASSN-14ae 11:08:40.11 +34:05:52.4 0.043671 2014-01-25 3.3 5 , 41 , 49 , 57 , 74 , 98 , 100 , 111 , 346 , 373 , 379
XMMSL1 J0740-85 07:40:08.19 -85:39:31.25 0.0173 2014-04-01 3.3 74 , 92 , 98 , 99 , 100 , 346
NGC 247 00:47:08.64 -20:45:36 0.000531 2014-07-01 1.0 23 , 64 , 74 , 98 , 100 , 346
CXOU J0332 03:32:38.8 -27:51:34 2.23 2014-10-01 0.0 109 , 346
DES14C1kia 03:34:47.49 -26:19:35 0.162 2014-11-11 0.0 5 , 58 , 74 , 98 , 100 , 346 , 373
ASASSN-14li 12:48:15.2253 +17:46:26.475 0.0206 2014-11-22 3.3 41 , 50 , 70 , 74 , 77 , 78 , 79 , 85 , 98 , 100 , 111 , 128 , 346 , 371 , 376 , 379

### v2

@@ V.4 Example 4: Validating Optical Transient Classifiers in the Rubin era
A sample application of the OTTER dataset is to apply a machine learning classifier to the available optical data to further test the ability of the classifier to recover the TDE classification from a sample of spectroscopically classified TDEs. We use the Finding Luminous and Exotic Extragalactic Transients (FLEET) algorithm, which is able to classify transients based on their optical light curves as SLSN, TDE, and “other” (Gomez et al., 2020, 2023) [^22] Note, while we report other classifications here, the classifications besides TDE and SLSN are not validated.. FLEET requires at least 4 detections in $g$-band or $r$-band, and is able to work with SDSS/ZTF filters (i.e. ugrizy) so we select all transients from the OTTER dataset (as of August 2025) that have enough photometry in these filters to model and run FLEET. This results in 83 transients, all of which are spectroscopically classified as TDEs.
@@
A final notable result is that FLEET appears to fail on the two TDEs in this sample that show on-axis jets (Sw J1644+57 and AT 2022cmc), returning P(TDE)=0% and finding very low probabilities for all classifications that FLEET is trained on. This is unsurprising given that these TDEs tend to appear much different photometrically from the non-jetted population (although they may share some similarities with the most luminous optical TDEs; Yao et al. 2025a). Additionally, given the extremely limited sample size of jetted TDEs, FLEET did not include any in their training sample. This may indicate that future TDE classifiers should attempt to classify jetted and non-jetted TDEs separately. In the future, we hope that OTTER will become a hub for curating training and validation datasets for future machine learning classifiers, particularly those that utilize multiwavelength datasets.
VI Conclusions
@@
For the initial release of this infrastructure, we also present the largest catalog of tidal disruption event candidates (TDE candidates). As of November 2025, the catalog includes from 240 TDE candidates that were classified in at least one published source. The API and web-application allow for easy access to this dataset, and the OTTER upload form will allow others to upload their datasets after publishing so that, as a field, we can continue to maintain this catalog.
We present four sample analyses using the available TDE data including a delay time analysis, a sub-population analysis, simple modeling, and validating a machine learning classifier. First, we perform a simple analysis of the TDE population and compute the delay time between the peak UV/Optical and X-ray emission to likely be $\lesssim 300$ days for the population. Second, we briefly discuss an analysis of the population of known Extreme Coronal Line Emitters (ECLEs) in the context of the broader TDE population. We recover the apparent relation that ECLEs appear to occur in dustier TDE hosts with a high statistical significance. Third, we model two TDEs with MOSFiT to show that it is very easy to apply models to this standard dataset. Fourth, and finally, we apply the FLEET transient classifier (Gomez et al., 2020, 2023) to the data in OTTER and find that FLEET successfully classifies $\sim 69\%$ of the sample as TDEs.
@@
We thank Edo Berger for useful discussion and feedback, particularly on the classification confidence flagging. We thank the staff at SciServer, namely Charles Glaser, Gerard Lemson, and Arik Mitschang, for their indispendable discussion and assistance deploying OTTER and making it accessible to the world. We thank Ryan Chornock, Iair Arcavi, Sara Faris, Margaret Shepherd, and Yvette Cendes for testing the OTTER web interface and providing feedback. We thank Megan Masterson and Muryel Guolo for taking the time to provide data from their respective population papers. We thank Mathieu Renzo for useful comments on the paper draft. We thank the anonymous referee for their insightful comments. N.F. acknowledges support from the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-2137419. KDA and CTC acknowledge support provided by the NSF through award SOSPA9-007 from the NRAO and award AST-2307668. KDA gratefully acknowledges support from the Alfred P. Sloan Foundation. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
Appendix A Metadata and References for each TDE Candidate
@@
The default name, coordinates, redshift and discovery date (i.e. the metadata) for all of the TDE candidates in OTTER is given in Table LABEL:tab:refs. The rightmost column gives all of the references associated with this transient event in OTTER, including metadata and photometry. The reference numbers are defined below LABEL:tab:refs.
Table 3: Default metadata information for the transients in OTTER
Transient Name RA [HMS] Dec [DMS] Redshift Discovery Date [MJD] C References
NGC 5905 15:15:23.32 +55:31:01.59 0.01124 1990-07-11 3.1 1 , 6 , 43 , 74 , 98 , 100 , 202 , 351
RX J1420+53 14:20:24.2 +53:34:11 0.147 1990-07-11 1.0 4 , 74 , 98 , 100 , 351
RX J1624+75 16:24:56.7 +75:54:57.5 0.0636 1990-08-25 1.0 3 , 43 , 74 , 98 , 100 , 202 , 351
RBS 1032 11:47:26.8 +49:42:59 0.026 1990-11-05 1.0 54 , 74 , 98 , 100 , 351
LEDA 095953 13:47:30.33 -32:54:50.63 0.0366 1991-07-06 1.0 11 , 74 , 98 , 100 , 351
RX J1242-11A 12:42:38.55 -11:19:20.8 0.05 1992-07-15 1.0 2 , 4 , 43 , 74 , 98 , 100 , 202 , 351
FIRST J153350.8+272729 15:33:50.88 +27:27:29.618 0.032430 1995-04-16 1.0 287
WINGS J1348 13:48:49.9 +26:35:57.49 0.062 1999-12-20 1.0 47 , 64 , 74 , 98 , 100 , 351
SDSSJ0159 01:59:57.64 +00:33:10.49 0.31167 2000-07-29 0.0 62 , 68 , 74 , 98 , 100 , 351
3XMM J152130.7+074916 15:21:30.75 +07:49:16.7 0.17901 2000-08-23 1.0 66 , 74 , 98 , 100 , 246 , 351
NGC 3599 11:15:26.95 +18:06:37.33 0.002699 2002-05-27 1.0 8 , 74 , 98 , 100 , 106 , 351
CSS100217 10:29:13 +40:42:20 0.147 2002-12-31 3.1 5 , 14 , 22 , 27 , 74 , 98 , 100 , 103 , 114 , 351
IGR J17361-4441 17:36:17.42 -44:44:05.98 0.000271 2003-03-22 1.0 17 , 56 , 74 , 98 , 100 , 351
2XMM J123103.2+110648 12:31:03.2 +11:06:48 0.11871 2003-07-13 0.0 36 , 37 , 46 , 74 , 98 , 110 , 351
SDSSJ0748 07:48:20.67 +47:12:14.23 0.0615 2003-10-23 3.1 25 , 34 , 74 , 98 , 100 , 351
SDSSJ1323 13:23:41.97 +48:27:01.26 0.08754 2003-12-02 1.0 8 , 43 , 74 , 98 , 100 , 106 , 202 , 351
D3-13 14:19:29.78806608 +52:52:06.3714576 0.3698 2004-01-15 3.3 7 , 9 , 21 , 42 , 74 , 98 , 100 , 111 , 202 , 351 , 391
SDSSJ1311 13:11:22.2 -01:23:45.2 0.18 2004-02-28 3.3 16 , 43 , 74 , 98 , 100 , 202 , 351
D1-9 02:25:16.96991014 -04:32:59.13393086 0.326 2004-06-22 3.3 7 , 9 , 42 , 74 , 98 , 100 , 202 , 351 , 391
2MASX J0249 02:49:17.28 -04:12:54 0.0186 2004-07-24 1.0 8 , 73 , 74 , 98 , 100 , 351
NGC 6021 15:57:30.72 +15:57:21.6 0.015607 2005-01-02 0.0 71 , 74 , 98 , 100 , 351
UGC 01791 02:19:53.52 +28:14:52.8 0.015881 2005-01-09 1.0 71 , 74 , 98 , 100 , 351
@@
PGC 133344 21:42:55.92 -30:07:58.8 0.02365 2005-03-17 1.0 71 , 74 , 98 , 100 , 351
3XMM J215022.4-055108 21:50:22 -05:51:08 0.055 2005-05-01 1.0 131 , 351
XMMSL1 J131952.3+225958 13:19:52 +22:59:58 0.023 2005-07-15 1.0 191 , 351
3XMM J150052.0+015452 15:00:52 +01:54:52 0.145 2005-07-23 1.0 74 , 98 , 115 , 351
PGC 170392 22:26:46.32 -15:01:22.8 0.016246 2005-11-28 1.0 71 , 74 , 98 , 100 , 351
XMMSL2 J144605.0+685735 14:46:05 +68:57:31 0.029 2006-08-22 3.3 150 , 202 , 351
TDE1 23:42:01.406 +01:06:29.3 0.136 2006-08-27 3.3 5 , 26 , 42 , 74 , 98 , 100 , 111 , 202 , 351 , 391
2XMMi J1847 18:47:25.16 -63:17:24.96 0.0353 2006-09-01 1.0 24 , 74 , 98 , 127 , 351
PGC 1127938 01:18:56.64 -01:03:10.8 0.02 2007-01-08 1.0 71 , 74 , 98 , 100 , 351
TDE2 23:23:48.615 -01:08:10.34 0.2515 2007-06-18 3.3 5 , 26 , 28 , 42 , 74 , 98 , 100 , 202 , 351 , 391
D23H-1 23:31:59.5344 +00:17:14.568 0.1855 2007-09-04 3.3 12 , 42 , 74 , 98 , 100 , 111 , 202 , 351 , 391
SDSSJ0952 09:52:09.56 +21:43:13.24 0.0789 2008-02-05 3.1 10 , 74 , 98 , 100 , 351
Dougie 12:08:47.87 +43:01:20.01 0.191 2009-01-17 3.1 61 , 74 , 98 , 100 , 351
J091225 09:12:25.01 +06:10:14.8 0.1453 2009-02-09 1.0 74 , 98 , 129 , 351
J100933 10:09:33.14 +23:22:55.8 0.0719 2009-02-20 1.0 74 , 98 , 129 , 351
J113527 11:35:27.29 +39:28:10.7 0.1084 2009-04-07 1.0 74 , 98 , 129 , 351
J133737 13:37:37.13 +20:23:51.7 0.0725 2009-04-12 1.0 74 , 98 , 129 , 351
J130819 13:08:19.12 +43:45:25.6 0.0365 2009-04-14 1.0 74 , 98 , 129 , 351
J141036 14:10:36.81 +26:54:25 0.1071 2009-04-24 1.0 74 , 98 , 129 , 351
J142401 14:24:01.66 +29:50:28.4 0.0855 2009-04-26 1.0 74 , 98 , 129 , 351
J145851 14:58:51.65 +17:50:57.9 0.1165 2009-04-29 1.0 74 , 98 , 129 , 351
J121116 12:11:16.47 +00:13:45.3 0.0763 2009-05-04 1.0 74 , 98 , 129 , 351
PTF09ge 14:57:03.1 +49:36:40.8 0.064 2009-05-07 3.3 5 , 13 , 41 , 55 , 74 , 98 , 100 , 111 , 351 , 391
SDSSJ1342 13:42:44.41 +05:30:56.2 0.0366 2009-05-17 1.0 34 , 45 , 74 , 98 , 100 , 129 , 351
J155223 15:52:23.29 +32:34:55.1 0.1277 2009-05-31 1.0 74 , 98 , 129 , 351
PTF09axc 14:53:13.06 +22:14:32.2 0.1146 2009-06-20 3.3 5 , 15 , 41 , 55 , 74 , 98 , 100 , 111 , 202 , 351 , 391
PTF09djl 16:33:55.94 +30:14:16.3 0.184 2009-07-24 3.3 5 , 15 , 41 , 55 , 74 , 98 , 100 , 111 , 113 , 351 , 391
J133837 13:38:37.6 +57:31:13.2 0.1269 2009-11-02 1.0 74 , 98 , 129 , 351
J030257 03:02:57.14 -08:29:42.9 0.1056 2009-11-08 1.0 74 , 98 , 129 , 351
PGC 1190358 15:05:28.56 +01:17:31.2 0.00766 2009-12-29 1.0 71 , 74 , 98 , 100 , 351
J123715 12:37:15.1 +60:12:07 0.2164 2010-01-23 1.0 74 , 98 , 129 , 351
PGC 1185375 15:03:50.4 +01:07:37.2 0.00523 2010-02-03 1.0 71 , 74 , 98 , 100 , 351
PGC 015259 04:29:21.84 -04:45:36 0.014665 2010-02-08 1.0 71 , 74 , 98 , 100 , 351
PS1-10jh 16:09:28.28 +53:40:23.99 0.1696 2010-05-16 3.3 5 , 40 , 42 , 67 , 74 , 98 , 100 , 111 , 202 , 351 , 391
SDSSJ1201 12:01:36.03 +30:03:05.52 0.146 2010-06-10 3.3 32 , 74 , 98 , 100 , 106 , 202 , 351
PTF10iya 14:38:41 +37:39:33.6 0.22405 2010-06-11 3.3 5 , 38 , 42 , 74 , 98 , 100 , 202 , 351
PTF10nuj 16:26:24.7 +54:42:21.6 0.132 2010-06-13 3.3 5 , 55 , 74 , 98 , 100 , 351
2022agi 01:02:50.007 -22:21:57.22 0.117835 2010-06-25 3.1 74 , 98 , 116 , 321 , 351
2MASX J01190869-3411305 01:19:08 -34:11:30 0.018 2010-07-14 1.0 162 , 351
2021zlg 20:42:44.74 +15:30:32.1 0.203 2010-08-15 3.3 74 , 98 , 114 , 282 , 351
UGC 03317 05:33:37.68 +73:43:26.4 0.004136 2010-09-22 1.0 71 , 74 , 98 , 100 , 351
PS1-11af 09:57:26.82 +03:14:00.924 0.4046 2010-12-30 3.3 5 , 52 , 74 , 98 , 100 , 202 , 351 , 391
IGR J12580+0134 12:58:05.1 +01:34:26 0.00411 2011-01-11 1.0 65 , 80 , 102 , 286
Sw J1644+57 16:44:49.9313 +57:34:59.6893 0.354 2011-03-28 3.3 29 , 30 , 31 , 33 , 39 , 44 , 74 , 76 , 81 , 98 , 100 , 126 , 234 , 351
Swift J2058.4+0516 20:58:19.9 +05:13:32 1.1853 2011-05-27 3.1 18 , 35 , 38 , 63 , 74 , 98 , 100 , 112 , 351
PTF11glr 16:54:06.13 +41:20:14.8 0.207 2011-05-28 3.3 5 , 55 , 74 , 98 , 100 , 351
XMMSL1 J063045.9-603110 06:30:46 -60:31:10 0.014 2011-12-01 1.0 20 , 72 , 93 , 351
Sw J1112-82 11:11:47.6 -82:38:44.44 0.89 2011-12-27 3.1 19 , 69 , 74 , 98 , 100 , 112 , 351
PS1-12yp 13:31:55.91 +23:54:05.7 0.581 2012-04-09 3.3 82 , 98 , 100 , 351
XMMSL1 J061927.1-655311 06:19:27 -65:53:07 0.0729 2012-11-12 3.1 48 , 351
PS1-13jw 08:44:53.56 +42:57:44.8 0.345 2013-02-06 3.3 74 , 98 , 114 , 351
ASASSN-14ae 11:08:40.11 +34:05:52.4 0.043671 2014-01-25 3.3 5 , 41 , 49 , 57 , 74 , 98 , 100 , 111 , 351 , 384 , 391
XMMSL1 J0740-85 07:40:08.19 -85:39:31.25 0.0173 2014-04-01 3.3 74 , 92 , 98 , 99 , 100 , 351
NGC 247 00:47:08.64 -20:45:36 0.000531 2014-07-01 1.0 23 , 64 , 74 , 98 , 100 , 351
CXOU J0332 03:32:38.8 -27:51:34 2.23 2014-10-01 0.0 109 , 351
DES14C1kia 03:34:47.49 -26:19:35 0.162 2014-11-11 0.0 5 , 58 , 74 , 98 , 100 , 351 , 384
ASASSN-14li 12:48:15.2253 +17:46:26.475 0.0206 2014-11-22 3.3 41 , 50 , 70 , 74 , 77 , 78 , 79 , 85 , 98 , 100 , 111 , 128 , 351 , 382 , 387 , 391